Tonight is a concert night. Three concerts – a classical music concert, a country music concert, and a rock music concert -- take place at the same time. You consider going to one of them. Suppose the ticket price of the classical concert is $80 and your willingness to pay is $100; the ticket price of the country music concert is $50 and your willingness to pay is $80; the ticket price of the rock music concert is $50 and your willingness to pay is $100. So you decide to go to the rock music concert. Apart from the ticket price, there is no other cost of seeing any of the three concerts. What is your ECONOMIC cost of attending the rock music concert?
Answer #1

Answer

here we compare the consumer surplus as the opportunity cost of the concert. as the opportunity cost is a benefit from the best alternative of the actions or time of activity or use of an asset or use of investment or money.

the consumer surplus =williness to pay -price

from rock music concert=100-50=50

from country music concert=80-50=30

from classical concert =100-80=20

so the consumer is going to get $50 instead of $30 as the best alternative for the surplus and that is the opportunity cost of the attending the rock music concert

The opportunity cost of attending the rock music concert is $30

the accounting cost =$50 which is the cost of the ticket

Economic cost =opportunity cost + accounting cost

=30+50

=$80

the economic cost is $80